OpenStudy (rayman5766):

What is the exact circumference of a circle with a radius of 12 inches? A. 6 inches B. 12 inches C. 24 inches D. 48 inches

OpenStudy (rayman5766):

C

OpenStudy (anonymous):

Neither of those appear to be ecen close to correct

OpenStudy (anonymous):

Circumference is calculated by two pi times radius, which in this case gives around 75

OpenStudy (rayman5766):

@Data_LG2

OpenStudy (anonymous):

\(\sf C= 2 \pi r\) is right :) you were given the radius and you are looking for circumference..plug them in \(\sf C= 2 \pi r =2(3.14)(12 in)= 75\ in\) hmmm.... are you sure you copied the right question?

OpenStudy (rayman5766):

huh yea I guess?

OpenStudy (rayman5766):

so c is correct

OpenStudy (anonymous):

how did you get C though?

OpenStudy (anonymous):

i'm getting 75 instead of 24

OpenStudy (rayman5766):

I guessed

OpenStudy (rayman5766):

those are the answer

OpenStudy (anonymous):

Since the Diameter of a circle with radius 12 is 24 I'll assume that's what the problem is? Am I right?

OpenStudy (rayman5766):

i guess

OpenStudy (anonymous):

"What is the exact circumference " -> that's what the problem is asking though ?

OpenStudy (rayman5766):

so my answer is 24 and that's how i get 75?

OpenStudy (rayman5766):

so its is c